OpenStudy (anonymous):

If the value of y is 7, what i the corresponding value of x? y=1/4x-1 pleasa show work

OpenStudy (mertsj):

First add 1 to both sides and get y+1=(1/4)x Now multiply both sides by 4 and get 4y+4=x Now replace y with 7 and find x. 32=x